OpenStudy (anonymous):

f''x = 2+cosx f(0)=-1 and f(pi/2)=0 Find f(x).

OpenStudy (anonymous):

find f(x)

OpenStudy (dumbcow):

integrate twice f'(x) = 2x +sin x + C f(x) = x^2 -cos x + Cx + K use the two initial values to find constants C,K

OpenStudy (anonymous):

but both given values are to plug in the original function...how can we solve for Cx?

OpenStudy (anonymous):

@dumbcow do u have any clue how to go further on?

OpenStudy (sidsiddhartha):

first put x=0 in the equation f(x) = x^2 -cos x + Cx + K u'll get K

OpenStudy (anonymous):

but how do we know which one we have to use first?

OpenStudy (sidsiddhartha):

if u put x=0 then c will be eliminated and u cal calculate K and then C but if u put x=pi/2 then both K and C will be there

OpenStudy (sidsiddhartha):

f(0)=-1 -1+k=-1 k=0 now u have f(x)= x^2 -cos x + Cx now put x=pi/2 f(pi/2)=0 0=pi^2/4+(C*pi)/2 c=-(pi/2)

OpenStudy (anonymous):

OMG i get it now!! thank uu

OpenStudy (sidsiddhartha):

np:)